OpenStudy (anonymous):

How do i put x^2+(k+3)x+4x=0 into the quadratic formula?

OpenStudy (lieutenantgeneral):

x^2 + 4x + 3 = 0 x = [-b ±√(b^2 - 4ac)]/2a a = 1 b = 4 c = 3 x = [-4 ±√(16 - 12)]/2 x = [-4 ±√4]/2 x = [-4 ±2]/2 x = [-4 + 2]/2 x = [-2]/2 x = -1 x = [-4 - 2]/2 x = [-6]/2 x = -3 ∴ x = -3 , -1

OpenStudy (anonymous):

@LieutenantGeneral, but where did k go? the whole question is to find the value of k, and i have to use the quadratic formula to find it... In the book it says, I have to use d=b^2 -4ac but doesn't seem right

hartnn (hartnn):

is it 4x in the end ?? or just 4 ?

OpenStudy (lieutenantgeneral):

www.wolframalpha.com/ or type it in on google

OpenStudy (anonymous):

it's 4x, @hartnn

hartnn (hartnn):

hmm, ok so first combine terms with x \(\Large (k+3)x +4x = (k+3+4)x = (k+7)x\) got this ?

OpenStudy (anonymous):

Ahh, I was confused on how to add the 4x to (k+3)x okay, got it!

hartnn (hartnn):

ok, now compare your equation with \(ax^2+bx+c=0\) can you get values for a,b,c first ?

OpenStudy (anonymous):

yeah, so a=1, b=k+7 and c=0 right?

hartnn (hartnn):

absolutely correct :) can you proceed ? ask if you still need help solving further.

OpenStudy (anonymous):

Yeah, i think I got it so far... I'll tag you, if i'm stuck... Thanks! :D

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

okay, i already need help... Is this right? d=(k+7)^2 -4(1)(0) =(k^2+14k)-0 -> d=k^2+14k?

OpenStudy (anonymous):

I just noticed i got the whole question wrong! It's 4k in the end, not 4x, @hartnn

hartnn (hartnn):

hahaha! i thought something was weird with the question :P

hartnn (hartnn):

find a,b,c values again :)

OpenStudy (anonymous):

I need to get glasses, lol :P okay so a=1, b=(k+3) and c=4k?

hartnn (hartnn):

correct!

OpenStudy (anonymous):

ahhh, see, that's what i have on my paper. but then i got confused because i wrote 4x instead of 4k.. lol so, d= (k+3)^2 -4(1)(4k) =k^2 +6k+9 - 16k ? =k^2 -10k + 9 ? do I factorise it and solve to get k? @hartnn

hartnn (hartnn):

you can factorise it but what do you actually want to find ?

OpenStudy (anonymous):

i need to find the values for k to be able to solve the equation

hartnn (hartnn):

you need values of "k" for which your quadratic equation has "EQUAL ROOTS" ?

hartnn (hartnn):

for equal roots, d= 0 k^2-10k +9 =0 and you get 2 integer values of k from here

OpenStudy (anonymous):

how do i get that? do i use the quadratic equation within the quadratic equation?

OpenStudy (anonymous):

i mean, like get two values for the square root to use in the quadratic equation ?

hartnn (hartnn):

so you just want to solve this : x^2+(k+3)x+4x=0 and get the answer in terms of k ?

OpenStudy (anonymous):

In order to solve x^2+(k+3)x+4x=0, i have to find the values of k first... or that's what it says

hartnn (hartnn):

\(b^2-4ac = k^2-10k+9\) \(\Large x_{1,2} = \dfrac{-(k+3)\pm \sqrt{k^2-10k+9}}{2}\) to get the value of "k", you really need more information than just that quadratic equation

OpenStudy (anonymous):

when i translate the question to english, it says this A6. Determine the values ​​of k that makes the equation has exactly one solution: I tried putting the equation into wolfamalpha.com and it shows the solutions for x : \[x = 1/2 (-\sqrt(k^2-10 k+9)-k-3)\] and/or \[x = 1/2 (\sqrt(k^2-10 k+9)-k-3)\] - but i don't know how, @hartnn

hartnn (hartnn):

YES! this is exactly what i needed!! "Determine the values ​​of k that makes the equation has exactly one solution"

hartnn (hartnn):

for a quadratic equation to have equal solutions, \(\Large b^2-4ac = 0 \)

hartnn (hartnn):

so, \(k^2 - 10k + 9=0\)

hartnn (hartnn):

so, \(k^2-9k-k+9=0\) can you solve further by factoring method ?

OpenStudy (anonymous):

Oh, so i do need to factorise it? i wrote that on the paper, hold on

hartnn (hartnn):

yes, you get good 2 integer values of k :)

OpenStudy (anonymous):

so, isn't it (k-1)(k-9)=0

hartnn (hartnn):

yes so what are the 2 values of k ?

OpenStudy (anonymous):

...i don't know, how do i find that? @hartnn

hartnn (hartnn):

\(\large AB=0 \implies A=0 ~or ~ B=0 \\ \large (k-1)(k-9)=0 \implies k-1=0 ~or~ k-9=0\)

OpenStudy (anonymous):

Okay, so is that k=1 or k=9 ?

hartnn (hartnn):

yes! thats it!

OpenStudy (anonymous):

Okay, so what do i do now? do i use the quadratic equation to find two solutions because the two different k values? @hartnn

hartnn (hartnn):

you have already completely solved your question, "Determine the values ​​of k that makes the equation has exactly one solution" they need values of k ! and you got them as k = 1 or 9

OpenStudy (anonymous):

ohhhhhhhhh, that was all? Well that easy once you knew the real question xP Sometimes it helps understanding the question first, haha... THANKS! :D Now I can hand in my assignment @hartnn

hartnn (hartnn):

you're welcome ^_^ good luck!

OpenStudy (anonymous):

:D Thanks, again!